I spent 2 fifths of my money. I now have $75.how much did I spend?
Answer by KMST(5328)

If you spent 2/5, you must have 3/5 left, because
1 - 2/5 = 3/5 .
If $75 is 3/5 of the money you had, 1/5 is $25=$75/3 .
Then, the 2/5 you spent is twice that, or $50.